# Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams

### v3

@@ 2.1 Experimental method
Multiple diagnostic methods were employed to investigate the flame oscillation dynamics, including: (1) a spatially and spectrally modulated multi-color laser absorption system for spatiotemporally resolved measurements of gas temperature and CO2 mole fraction; (2) an image-intensified high-speed camera for spatiotemporally resolved imaging of the line-of-sight integrated OH* chemiluminescence (as a proxy for the local heat release rate); and (3) a photomultiplier tube (PMT) for microsecond-resolved measurements of global chemiluminescence variations and transient phenomena.
Figure 1: The current experimental setup: (a) schematic of the overall diagnostic system and burner; (b) detailed configuration of the porous-plug burner; (c) example of a burner-stabilized flame; (d) detailed view of the water cooling channel; (e) photograph of the static mixer.
@@ 2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
For a pulsating flame, the frequency content of the total chemiluminescence signal was first analyzed using fast Fourier transform (FFT). The results were expressed in terms of normalized power spectral density (NPSD), as shown in Eqn. (1).
@@
A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated. Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
@@
This optimization problem was solved numerically using a brute-force search algorithm, and to improve computational efficiency and robustness, the number of harmonics was truncated to M = 5. In the current study, this method was able to determine the primary oscillation frequency with a typical uncertainty of less than 0.1 Hz (defined by the frequency offset at 95% maximum total power), which was an order of magnitude better than the conventional FFT.
Once the primary oscillation frequency was determined, the chemiluminescence images of the pulsating flames were further analyzed using phase-locked cycle averaging. The recorded image sequence was divided into sub-sequences corresponding to individual cycles of oscillation, and images captured at the same relative time within each cycle were averaged across different sub-sequences. The starting frame of each sub-sequence was fine-tuned through cross-correlation analysis, using the spatially integrated chemiluminescence signal as a reference, which effectively removed random jitters in the oscillation cycles. This procedure was executed iteratively until convergence was achieved, usually within 2-3 iterations. This phase-locked cycle averaging successfully removed random fluctuations and revealed the underlying physical trends consistent across cycles, without degrading the effective temporal resolution of the current measurements (100 $\mu s$).
@@ 2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
Spatially resolved measurements of local gas temperature were achieved by scanning the laser beams across the flame. An example measurement of a pulsating CH4-O2 flame at lower flow rates ($\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM, respectively) is illustrated in Fig. 2. Under the specific conditions of this measurement, the flame exhibited a regular mode of oscillation that was periodic in time and nearly axisymmetric in space, as confirmed by time-synchronized chemiluminescence imaging. For oscillating flames exhibiting axial symmetry, a reconstruction scheme using constrained optimization was developed to infer the spatial distribution of temperature from the absorbance data at different beam locations.
@@
The line-of-sight absorbance $\alpha(t;y_{i})$ was determined at each beam location $y_{i}$ from the transmitted laser intensity $I(t;y_{i})$ and a reference intensity profile $I_{0}(t)$ measured in the absence of absorption, based on the Beer-Lambert relation:
$\alpha(t;y_{i})$ $=-\ln\big{[}I(t;y_{i})/I_{0}(t)\big{]}$ (4)
From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment. A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
$A_{k}(y_{i})=\int_{\nu_{k-}}^{\nu_{k+}}\alpha(\nu;y_{i})d\nu$ (5)
To determine the spatial distribution of gas temperature, the entire spatial domain of interest was discretized using a number of sample points $r_{s}$ along the radial direction. These sample points were selected based on the effective spatial resolution of the current laser diagnostic (1 mm, as defined by the laser beam diameter). At each sample point, a local gas temperature $T_{s}$ and a CO2 mole fraction $X_{s}$ were assigned. The integrated absorbance of each absorption feature was evaluated first at each sample point $A^{0}_{k}(r_{s};T_{s},X_{s})$ and then summed along each line of sight (see Fig. 2(b)), yielding:
$A_{k}^{0}(r_{s};T_{s},X_{s})$ $=PX_{s}S_{k}^{0}(T_{s})$ (6)
and
$A_{k}(y_{i})$ $=\int_{-\infty}^{+\infty}A^{0}_{k}\Big{(}\sqrt{x^{2}+y_{i}^{2}}\Big{)}dx\approx\sum_{s=1}^{\#pts.}w_{is}\cdot A_{k}^{0}(r_{s};T_{s},X_{s})$ (7)
In the above equations, a constant flame pressure of $P$ = 1 atm was assumed, and the linestrength of each transition feature, $S_{k}^{0}$, was calculated as a function of the local gas temperature from the HITEMP spectroscopy model [46]. The line-of-sight integration was numerically evaluated using Gauss–Legendre quadrature, with linear interpolation between sample points. This approach effectively approximated the integral as a weighted sum of the integrated absorbances at selected sample points. The weight factor $w_{is}$ was a geometric constant independent of the laser wavenumber $\nu$ and the local gas properties $T_{s}$ and $X_{s}$. The integrated absorbances were assembled into two column vectors of $\boldsymbol{A^{0}(T,X)}=\{A^{0}_{k}(r_{s};T_{s},X_{s})\}$ and $\boldsymbol{A}=\{A_{k}(y_{i})\}$, with $\boldsymbol{T}$ and $\boldsymbol{X}$ representing $\{T_{i}\}$ and $\{X_{i}\}$, respectively. All weight factors $\{w_{is}\}$ were compiled into a constant matrix $\boldsymbol{W}$. The optimal values of $\boldsymbol{T}$ and $\boldsymbol{X}$ were determined by solving the following nonlinear optimization problem:
$(\boldsymbol{T^{*},X^{*}})={\operatorname{argmin}}\big{\|}\boldsymbol{A-WA^{0}(T,X)}\big{\|}_{2}^{2}+\lambda_{T}^{2}\big{\|}\boldsymbol{LT}\big{\|}_{2}^{2}+\lambda_{X}^{2}\big{\|}\boldsymbol{LX}\big{\|}_{2}^{2}$ (8)
The last two terms in Eqn. 8 represent smoothness regularization, with $\boldsymbol{L}$ being the discrete Laplacian matrix and $\lambda_{T}$ and $\lambda_{X}$ serving as the regularization strength factors. The values of $\lambda_{T}$ and $\lambda_{X}$ were generally kept small to avoid blurring the reconstructed distributions, while being sufficiently large to ensure numerical stability and suppress random fluctuations; in the current study, they were set to $1\times 10^{-5}$ and $1\times 10^{-1}$, respectively. Eqn. 8 was solved iteratively using the sequential quadratic programming method. The resulting spatial distributions of temperature and CO2 mole fraction are shown in Fig. 4(c), with a typical measurement precision of 1%. Further discussion on the uncertainty analysis, as well as the choice of $\lambda_{T}$ and $\lambda_{X}$, can be found in a recent work by the authors [44].
@@ 3.1 Pulsation frequencies
Figure 3: (a) Representative OH* NPSD spectra of pulsating flames at three different conditions. Top (blue): at $\phi$ = 1.72, $\eta$ = 1.00, the NPSD spectrum consists of equally spaced harmonics of a single fundamental frequency, indicating the presence of single-mode oscillation. Middle (green): at $\phi$ = 1.72, $\eta$ = 0.71, low-frequency flickering instability appears, which couples with the single-mode oscillation and creates sidebands around the harmonics. Bottom (red): at $\phi$ = 1.47, $\eta$ = 0.71, multiple fundamental oscillation frequencies (and their harmonics) are observed. $f_{\text{OSC}}$: the primary oscillation frequency; $f_{\text{FI}}$: the flickering instability frequency. (b) Total harmonic power analysis of the three pulsating flames.
@@
At higher flow velocities, a second, low-frequency oscillation appeared around 10 Hz and created sidebands around the primary oscillation frequency, as shown in the middle panel of the figure. Across the experimental conditions explored in the current study, the frequency of this second oscillation was found to be relatively stable (9 - 12 Hz) and largely independent of the equivalence ratio and dilution level, suggesting that it must have originated from a mechanism different from the high-frequency thermo-diffusive pulsation.
@@
To further investigate the nature of this low-frequency oscillation, additional experiments were conducted using a CMOS camera to monitor the chemiluminescence signal of an undiluted, moderately fuel-rich flame from the side, as illustrated in Fig. 4. Under these conditions, the flame did not exhibit any high-frequency pulsation; only a low-frequency oscillation at a fundamental frequency of 11 Hz was observed. The results indicate that the low-frequency component corresponds to flickering-type flame dynamics that is distinctively different from flame front pulsation. For example, throughout a flame flickering cycle, the base of the flame remained largely stationary, whereas the tail of hot burnt gas downstream of the flame exhibited an oscillating motion of elongation and contraction along the axial direction.
It turns out that the current observation of this rich premixed flame agrees well with classic studies on diffusion flames, in terms of both the overall phenomenon and the range of oscillation frequencies (e.g., [47, 48]). According to these studies, such low-frequency oscillations of the flame result from buoyancy-induced flow surrounding the flame. The velocity difference between the hot burnt gases and the cold ambient air creates an unstable shear layer, which rolls up into toroidal vortices and sheds periodically. The current study thus refers to this low-frequency oscillation as the flame flickering instability, with its frequency denoted as $f_{\rm FI}$.
@@
Figure 4: representative example on the low-frequency oscillation of a methane-oxygen flame at $\phi$ = 1.20, $\eta$ = 1.00. (a) the FFT spectrum of CH* normalized power spectral density (NPSD), (b) sequential side-view images of the flame.
@@ 3.2 Single-mode oscillation
It is worth noting that, although the measured images were line-of-sight integrated, no significant cancellation effect was expected in the top-view images because the flame front remained close to a geometrically thin sheet throughout the oscillation cycle. Besides, the flame motion occurred primarily in the vertical direction, as revealed by additional side-view measurement of a single-mode pulsating flame shown in Fig. 6.
@@
Fig. 5(c) illustrates the dynamic evolution of temperature distribution at a height above the burner (HAB) of 2.5 mm. A similar ring-type oscillation can be observed in the temperature distribution, but with a phase lag compared to the chemiluminescence signal. At the specific conditions of this experiment, the maximum local temperature is found to be approximately 2000 K, significantly lower than the adiabatic flame temperature ($\approx$ 3000 K) due to heat loss at the water-cooled burner surface.
@@
Fig. 5(d) shows the evolution of line-of-sight-averaged temperature within one oscillation cycle. In the current study, a high level of measurement precision (on the order of $\pm$ 5 K) has been achieved through cycle averaging, thereby allowing small variations to be discerned. The average temperature is seen to oscillate between 1760 and 1790 K, with a phase lag of 100 - 150 degrees relative to the chemiluminescence signal. The overall 1-$\sigma$ uncertainties in these measurements are estimated to be 2 %, based on the uncertainty contributions from the spectral fitting (less than 0.1%), spectroscopy model (less than 1%), and small fluctuations in the laser intensity baseline (less than 1%).
Fig. 5(e) presents the vertical distribution of average temperature along the burner diameter, measured at different heights above the burner (HAB) and at various times during a single oscillation cycle. Its dynamic evolution reveals an interesting phenomenon that cannot be captured by the chemiluminescence imaging alone, namely, the existence of higher and lower average temperature regions and their vertical motion with time, which implies alternating puffs of hot and cool gas traveling in the vertical direction. Further investigation into the 2D/3D structure of the traveling flame fronts, especially in conjunction with velocity field measurement, may warrant future studies.
@@
Figure 5: Spatiotemporally resolved measurements of a single-mode pulsating flame ($\phi$ = 1.33, $\eta$ = 1.00) at different times in the oscillation cycle. (a) Spatially integrated chemiluminescence signals from the high-speed camera (black) and the PMT (blue), normalized and phase-locked averaged over 90 cycles. (b) Phase-locked averaged chemiluminescence images. (c) Dynamic evolution of the gas temperature distribution at HAB = 2.5 mm. (d) Evolution of the average temperature along the burner diameter. Error bars denote the standard deviation of the measurements. (e) Evolution of the vertical distribution of average gas temperature along the burner diameter within a single oscillation cycle.
Figure 6: Side-view chemiluminescence images of a single-mode pulsating flame at $\phi$ = 2.00, $\eta$ = 1.00.
@@
Figure 7: Chemiluminescence signal illustrating the spatiotemporal modes of flame oscillation. Selected frames of the cycle-averaged chemiluminescence images, after the subtraction of mean intensity, are presented in pseudocolors, with red and blue representing high and low signals, respectively. (a) Single-mode oscillation influenced by flickering instabilities at $\phi$ = 1.72, $\eta$ = 0.71. The dominant flame oscillation mode appears to be a fast centripetal traveling wave that is nearly axisymmetric. The flickering instabilities are manifested as a slow edge-mode perturbation that modulates the relative phase of the fast flame oscillation without significantly affecting its amplitude and spatial distribution. (b) Multi-mode oscillation at $\phi$ = 1.47, $\eta$ = 0.71. The integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies that interfere with each other. The propagation directions and spatial patterns of flame oscillations during periods of constructive and destructive interference are drastically different.

### v4

@@ 2.1 Experimental method
Multiple diagnostic methods were employed to investigate the flame oscillation dynamics, including: (1) a spatially and spectrally modulated multi-color laser absorption system for spatiotemporally resolved measurements of gas temperature (10 kHz, 2% accuracy); (2) an image-intensified high-speed camera for spatiotemporally resolved imaging of the line-of-sight integrated OH* chemiluminescence (10 kHz, as a proxy for the local heat release rate); (3) a photomultiplier tube (PMT) for microsecond-resolved measurements of global chemiluminescence variations and transient phenomena; (4) a Z-type high-speed schlieren system and (5) a pulsed OH-PLIF diagnostic for side-view measurements of the flames.
Figure 1: The current experimental setup. (a) Configuration of the porous-plug burner. (b) Example of a burner-stabilized flame. (c) Detailed view of the water cooling channel. (d) Photograph of the static mixer. (e) Schematic of the flame chemiluminescence and the tunable diode laser diagnostics. (f) The side-view schlieren diagnostic. (g) The side-view OH-PLIF diagnostic system.
@@
Side-view schlieren imaging of the flames was conducted using a Z-type configuration with a pair of concave mirrors, a 5-W LED, and a vertical knife edge located at the focal plane of the downstream mirror. The knife edge was slightly offset from the focus to enhance image contrast, enabling clear observation of minute density variations in the flow field during the experiment. Schlieren images were recorded with the Phantom v611 high-speed camera at a frame rate of 1 kHz.
OH-PLIF measurements were performed using a nanosecond-pulsed tunable dye laser (LiOPStar-N, LIOP-TEC) at a repetition rate of 10 Hz. The laser wavelength was tuned to 282.997 nm, targeting the P(1.5) + Q(1.5) + R(2.5) transition cluster of the OH A–X (1,0) band, and the pulse energy was measured to be approximately 30 mJ. The laser beam was expanded into a horizontal light sheet using a set of cylindrical lenses, achieving a $2\sigma$ thickness of less than 0.3 mm across the burner surface. The OH-PLIF signal was collected with the same intensified camera used for OH* chemiluminescence, and spectrally filtered between 300–320 nm to suppress background flame emission. The camera was gated (with an exposure time of 10 $\mu s$) and synchronized to the laser pulses. During each OH-PLIF measurement, the PMT signal of OH* chemiluminescence was recorded simultaneously. The superposition of laser-induced fluorescence and laser scattering on OH* chemiluminescence produced local spikes in the PMT signal, enabling determination of the relative phase of each OH-PLIF image within the flame oscillation cycle.
@@ 2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
The frequency content of the total chemiluminescence signal was first analyzed using fast Fourier transform (FFT). The results were expressed in terms of normalized power spectral density (NPSD), as shown in Eqn. (1).
@@
The fundamental frequency of oscillation, $f_{osc}$, was extracted using a harmonic power analysis method developed in a previous study of the authors [52]. Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
@@
This optimization problem was solved numerically using a brute-force search algorithm, and to improve computational efficiency and robustness, the number of harmonics was truncated to M = 5. In the current study, this method was able to determine the primary oscillation frequency with a typical uncertainty of less than 0.1 Hz (defined by the frequency offset at 95% maximum total power). Once the primary oscillation frequency was determined, the chemiluminescence images of the pulsating flames were further analyzed using phase-locked cycle averaging, following the same procedure as explained in [52].
@@ 2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
Spatially resolved measurements of local gas temperature were achieved by scanning the laser beams across the flame. An example measurement of a pulsating CH4-O2 flame at lower flow rates ($\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM, respectively) is illustrated in Fig. 2. Under the specific conditions of this measurement, the flame exhibited a regular mode of oscillation that was periodic in time and nearly axisymmetric in space, as confirmed by time-synchronized chemiluminescence imaging. For oscillating flames exhibiting axial symmetry, a reconstruction scheme using constrained optimization was developed to infer the spatial distribution of temperature from the absorbance data at different beam locations. Further details of the reconstruction scheme can be found in a previous study by the authors [58] and are therefore omitted here for brevity.
@@ 3.1 Pulsation frequencies
Figure 3: (a) Representative OH* NPSD spectra of pulsating flames at three different conditions. Top (blue): $\phi$ = 1.72, $\eta$ = 1.00; middle (green): $\phi$ = 1.72, $\eta$ = 0.71; bottom (red): $\phi$ = 1.47, $\eta$ = 0.71. $f_{\text{OSC}}$: the primary oscillation frequency; $f_{\text{FI}}$: the flickering instability frequency. (b) Total harmonic power analysis of the three pulsating flames.
@@
At higher flow velocities, a second, low-frequency oscillation appeared around 10 Hz and created sidebands around the primary oscillation frequency, as shown in the middle panel of the figure. Across the experimental conditions explored in the current study, the frequency of this second oscillation was found to be relatively stable (mostly between 9 and 12 Hz) and largely independent of the equivalence ratio and dilution level, suggesting that it must have originated from a mechanism different from the high-frequency thermo-diffusive pulsation.
@@
An example set of measurements for such low-frequency oscillation is shown in Fig. 4. These measurements were conducted on an undiluted, moderately fuel-rich flame of equivalence ratio $\phi$ = 1.20. Under such conditions, the flame did not exhibit any high-frequency pulsation; only a low-frequency oscillation at a fundamental frequency of 11 Hz was observed. The results indicate that the low-frequency component corresponds to flickering-type flame dynamics that is distinctively different from flame front pulsation. For example, throughout a flame flickering cycle, the base of the flame remained largely stationary, whereas the tail of hot burnt gas downstream of the flame exhibited an oscillating motion of elongation and contraction along the axial direction.
@@
Figure 4: Example low-frequency oscillation of a methane-oxygen flame at $\phi$ = 1.20, $\eta$ = 1.00. (a) the FFT spectrum of OH* normalized power spectral density (NPSD), (b) sequential side-view images of the flame, (c) side-view schlieren images of the toroidal vortex formation and shredding.
These observed low-frequency oscillations are consistent with the general frequency range of buoyancy-controlled flickering of diffusion flames reported in the early study by Buckmaster and Peters [7]. The general dynamics of these oscillations agree qualitatively with the results of the planar visualization study by Chen et al. [12], in which the flickering instability is correlated with the periodic shredding of large toroidal vortices generated by buoyancy‑driven Kelvin–Helmholtz instability outside the luminous flame. The vortex-dynamical picture of flame flickering instability has been revisited by a multitude of studies under various fuel types, equivalence ratios, burner diameters, and gravity levels (for example, [44, 43, 54, 55]). A consensus has emerged that flickering is primarily caused by buoyant flow instability and toroidal vortices, at least at sufficiently small Reynolds numbers, and is distinct from flame extinction and re-ignition or direct modulation of the flame’s chemical reaction/heat release rate. In this sense, flickering instability is not restricted to diffusion flames but applies to premixed flames as well. The present study thus refers to this low-frequency oscillation as flame flickering instability, with its frequency denoted as $f_{\rm FI}$.
Previous studies on the flickering frequency of buoyant diffusion flames (for example, [50, 4, 17, 11, 32, 21]) have suggested the Strouhal number ($St=f_{FI}D/U$, where $D$ is the burner diameter and $U$ is the unburnt gas velocity) correlates with the Froude number ($Fr=gD/U^{2}$) as $St\propto Fr^{-0.5}$. In this study, a total of 224 measurements were conducted on the oscillation frequencies of flickering flames without thermo-diffusive pulsations. The resulting non-dimensional frequency, expressed in terms of St, is plotted as a function of the inverse Froude number in Fig. 5. The best numerical fit to the current data yields a scaling relation of $St\propto Fr^{-0.44}$, which agrees with the previous scaling of buoyancy-controlled flickering.
Figure 5: Scaling of the measured non-dimensional flame flickering frequency (St) versus the inverse Froude number (1/Fr). Results are shown for purely flickering flames without thermo-diffusive pulsations.
The present study also examined several other mechanisms that could potentially caused the low-frequency flame oscillations, including: (a) large-scale hydrodynamic shedding, (b) acoustic coupling, and (c) upstream forcing. None of these were found to be the actual cause. For hydrodynamic vortex shedding, the observed Strouhal number ($St$) varies significantly across different flame conditions, which differs markedly from the classical, nearly constant $St$ associated with vortex shedding from bluff bodies. Regarding acoustic coupling, the resonant acoustic frequency of the burner, estimated from the speed of sound and the effective cavity length inside the burner, is approximately 1250 Hz – two orders of magnitude higher than the observed flickering frequency. Upstream forcing was minimized by precisely controlling and continuously monitoring the mass flow rates throughout the experiments, and no upstream oscillation near 10 Hz was observed.
@@ 3.2 Single-mode oscillation
It is worth noting that, although the measured chemiluminescence images were line-of-sight integrated, no significant cancellation effect was expected in the top-view images because the flame front remained close to a geometrically thin sheet throughout the oscillation cycle. This was confirmed by side-view measurements of a single-mode pulsating flame, the results of which are shown in Fig. 6(c).
The dynamic evolution of the flame structure can also be visualized from OH-PLIF images, as shown in Fig. 6(d). The PLIF measurements revealed additional details of the flame structure downstream of the flame front, which were obscured in the chemiluminescence imaging by the rapid quenching of electronically excited OH radicals (OH*) to their ground state. From the OH-PLIF measurements, an interesting phenomenon was observed: alternating regions of higher and lower OH signal were moving both horizontally and vertically within an oscillation cycle, suggesting that the flame oscillation mode followed a traveling-wave pattern. This was also confirmed by flame temperature measurements discussed below.
@@
The evolution of temperature distribution at a height above the burner (HAB) of 2.5 mm is illustrated in Fig. 6(e). A similar ring-type oscillation can be observed. At the specific conditions of this experiment, the maximum local temperature is found to be approximately 2000 K, significantly lower than the adiabatic flame temperature ($\approx$ 3000 K) due to heat loss at the water-cooled burner surface.
@@
Fig. 6(f) shows the evolution of line-of-sight-averaged temperature within one oscillation cycle. In the current study, a high level of measurement precision (on the order of $\pm$ 5 K) has been achieved through cycle averaging, thereby allowing small variations to be discerned. The average temperature is seen to oscillate between 1760 and 1790 K, with a phase delay of 100 - 150 degrees relative to the chemiluminescence signal. The overall 1-$\sigma$ uncertainties in these measurements are estimated to be 2 %, based on the uncertainty contributions from the spectral fitting (less than 0.1%), spectroscopy model (less than 1%), and small fluctuations in the laser intensity baseline (less than 1%).
@@
Figure 6: Spatiotemporally resolved measurements of a single-mode pulsating flame ($\phi$ = 1.33, $\eta$ = 1.00) at different times in its oscillation cycle. (a) Spatially integrated chemiluminescence signals from the high-speed camera (black) and the PMT (blue), normalized and phase-locked averaged over 90 cycles. (b)-(c) Phase-locked averaged images of top-view and side-view chemiluminescence, respectively. (d) Instantaneous images of side-view OH-PLIF. (e) Dynamic evolution of the gas temperature distribution at HAB = 2.5 mm. (f) Evolution of the average temperature along the burner diameter. Error bars denote the standard deviation of the measurements. (g) Evolution of the vertical distribution of average gas temperature along the burner diameter within a single oscillation cycle.
@@
Figure 7: Chemiluminescence signal illustrating the spatiotemporal modes of flame oscillation. Selected frames of the cycle-averaged chemiluminescence images, after the subtraction of mean intensity, are presented in pseudocolors, with red and blue representing high and low signals, respectively. (a) Single-mode oscillation influenced by flickering instabilities at $\phi$ = 1.72, $\eta$ = 0.71. (b) Multi-mode oscillation at $\phi$ = 1.47, $\eta$ = 0.71.
The vertical distribution of average temperature along the burner diameter, measured at different heights above the burner (HAB) and at various times during a single oscillation cycle, is presented in Fig. 6(g). The existence of higher and lower average-temperature regions and their vertical motion with time correlates with the PLIF signal shown in Fig. 6(d), indicating alternating puffs of hotter and cooler gases traveling in the vertical direction, with a phase delay similar to that of Fig. 6(f). Such a phase delay suggests that the flame oscillation is not caused by heat loss to the burner surface alone, and there must be an additional mechanism – such as the differential transport between mass and heat – that interacts with the heat loss and jointly drives the oscillation cycle. Further investigation into the 2D/3D structure of the traveling flame fronts and the detailed mechanism of flame oscillation, especially in conjunction with velocity-field measurement, is warranted for future studies.